Jake is 3 Years older Than kyla and 2 Years younger than larry. The total age is 41 .Find jake's age
Found 2 solutions by macston, addingup:
Answer by macston(5194)

.
J=Jake's age; K=Kyla's age=J-3; L=Larry's age=J+2
.
J+K+L=41
J+(J-3)+(J+2)=41
3J-1=41
3J=42
J=14
.
ANSWER: Jake is 14 years old.
.

Answer by addingup(3677)

K= J-3
L= J+2
-----------------------------
J-3+J+J+2 = 41
J-1+J+J = 41
3J = 42
J = 14 this is how old Jake is.
Kyla is 3 years younger: 14-3= 11
Larry is 2 years older: 14+2= 16
J